# Form and meaning co-determine the realization of tone in Taiwan Mandarin spontaneous speech: the case of T2-T3 and T3-T3 tone sandhi

### v1

@@ Abstract
In Standard Chinese, Tone 3 (the dipping tone) becomes Tone 2 (rising tone) when followed by another Tone 3. Previous studies have noted that this sandhi process may be incomplete, in the sense that the assimilated Tone 3 is still distinct from a true Tone 2. While Mandarin Tone 3 sandhi is widely studied using carefully controlled laboratory speech (Xu, 1997) and more formal registers of Beijing Mandarin (Yuan and Chen, 2014), less is known about its realization in spontaneous speech, and about the effect of contextual factors on tonal realization. The present study investigates the pitch contours of two-character words with T2-T3 and T3-T3 tone patterns in spontaneous Taiwan Mandarin conversations. Our analysis makes use of the Generative Additive Mixed Model (GAMM, Wood, 2017) to examine fundamental frequency (f0) contours as a function of normalized time. We consider various factors known to influence pitch contours, including gender, speaking rate, speaker, neighboring tones, word position, bigram probability, and also novel predictors, word and word sense (Chuang et al., 2024). Our analyses revealed that in spontaneous Taiwan Mandarin, T3-T3 words become indistinguishable from T2-T3 words, indicating complete sandhi, once the strong effect of word (or word sense) is taken into account. For our data, the shape of f0 contours is not co-determined by word frequency. In contrast, the effect of word meaning on f0 contours is robust, as strong as the effect of adjacent tones, and is present for both T2-T3 and T3-T3 words.
@@
Keywords: Tone 3 sandhi, tonal assimilation, GAMM, word-specific tonal realization, frequency effect
@@
There is an increasing interest in spontaneous speech in phonetics, psycholinguistics, and related fields. While carefully controlled speech from laboratory experiments can provide valuable insights into specific aspects of speech production, spontaneous speech allows researchers to study authentic language use in everyday conversations. Spontaneous speech differs from laboratory speech in various aspects, such as the choice of words, syntactic patterns, tones and intonation, and phonological assimilation (Tucker and Ernestus, 2016). For instance, the pronunciation of Mandarin tones in spontaneous speech in everyday use differs from how it is described in textbooks and second-language learning classrooms, posing a huge challenge for second language learners of Mandarin.
@@
Mandarin Chinese is a tonal language, with each syllable being distinguished by a specific lexical tone in addition to vowels and consonants. This tonal differentiation contributes to distinguishing between words’ meanings. The main acoustic correlate of Mandarin tones is the shape of words’ fundamental frequency (f0) contours. Mandarin Chinese has four lexical tones: a high-level tone (T1), a rising tone (T2), a dipping tone (T3), and a falling tone (T4), along with a neutral or floating tone (Chao, 1968). Chao (1968) describes the neutral tone as being unstressed, noticeably weaker in intensity, and shorter in duration. When a syllable is placed in a disyllabic word or connected speech, tone sandhi takes place. Tone sandhi refers to the modulation of a given syllable’s lexical tones by the tones of other syllables in its context. Tone 3 sandhi is the most widely studied tone sandhi phenomenon in Mandarin Chinese. In a disyllabic word with the tone pattern T3-T3, the first T3 is typically pronounced as T2, leading to the perception that the T3-T3 tone pattern sounds similar to the T2-T3 tone pattern. There has been a long-standing debate about whether this tone sandhi results in incomplete or complete assimilation of initial T3 to a initial T2 (Yuan and Chen, 2014).
@@
Although Tone 3 sandhi has long been studied for laboratory speech (Shih, 1986; Xu, 1997), its realization in spontaneous speech has been less extensively investigated. It has been observed that in general, several standard acoustic cues for individual tones, such as f0, are greatly influenced by contextual factors or even reduced to a large degree in spontaneous speech, rendering them unreliable (Brenner, 2013). The majority of research in Mandarin Tone 3 sandhi is based on words that are articulated in isolation or words embedded in simple sentences or careful speech elicited in the lab using experimental tasks (Warner, 2011; Wagner et al., 2015). In order to gain a more comprehensive understanding of the production of Tone 3 sandhi, it is essential to examine it in the speech style that speakers and listeners most frequently and commonly use in everyday settings. The study of Yuan and Chen (2014), however, addresses the realization of Tone 3 sandhi as it is realized in spontaneous spoken Mandarin, using corpora of standard Mandarin telephone speech and Mandarin broadcast news speech. The present study aims to investigate the realization of disyllabic words with T2-T3 and T3-T3 tone patterns in spontaneous spoken Taiwan Mandarin in face-to-face conversations. In what follows, section 1 provides further details on what is currently known about Tone 3 sandhi. Section 2 introduces our materials and the methods we used for data analysis. Results are presented in section 3. Our study concludes with a summary and discussion.
@@ 1 Incomplete neutralization of Tone 3 sandhi
Most studies have reported incomplete neutralization for Tone 3 sandhi in Mandarin Chinese. The Sandhi Rising tone (SR tone, the first T3 in T3-T3) is found to exhibit lower pitch compared to the Lexical Rising tone (LR tone, T2 in T2-T3), with mean differences ranging from 3.2 Hz to 20 Hz (Cui et al., 2020; Kratochvil, 1984; Myers and Tsay, 2003; Shen, 1990; Xu, 1993, 1997; Zee, 1980). The SR tone has also been described as tending to have less f0 excursion in the pitch contour across various varieties of Mandarin. For instance, Xu (1997) examined disyllabic non-word /ma-ma/ sequences with 16 possible Mandarin bi-tonal combinations embedded in carrier sentences produced by Beijing Mandarin native speakers. This study revealed that although both exhibited similar fall-rise-fall f0 contours, T3-T3 has a slightly lower pitch than T2-T3 throughout both syllables. This observation suggests for laboratory speech, the SR tone does not fully assimilate with the LR tone. However, most studies addressing Tone 3 sandhi in Taiwan Mandarin have reported that the SR tone and the LR tones are more similar to each other than in other varieties such as Beijing Mandarin. These differences in realization may remain visible but are not well supported from statistical analysis (Cheng et al., 2013; Fon and Chiang, 1999; Myers and Tsay, 2003; Peng, 2000).
@@
Only a few studies have examined Tone 3 sandhi in connected or spontaneous speech. An early study by Kratochvil (1987) analyzed the speech of a speaker of Beijing Mandarin, who was asked to generate sentences using words presented in list format. Using discriminant analysis as classifier, he found that most of the tokens with the sandhi T3 were grouped with the lexical T3 rather than the lexical T2. This study concluded that speakers of Mandarin are adjusting the underlying T3, but the resulting SR tone is yet not equivalent to a T2.
@@
More recently, Yuan and Chen (2014) analyzed the realization of Tone 3 sandhi in Mandarin bi-character words in connected speech from a large corpus of telephone conversations and formal news broadcasts. They found that, despite the remarkable similarity, the LR tone displays higher and longer f0 rising than the SR tone. This low-level acoustic difference in the magnitude and the time period of f0 rise is consistent with previous results obtained using carefully controlled speech. Following up on the study by Xu (1997), Wu et al. (2021) analyzed Mandarin disyllabic words with 16 tonal combinations, but used large corpora of journalistic speech. They found that the SR tone in the T3-T3 sequence is often realized as other tones, suggesting that the tone sandhi rule in connected speech is more likely to be a tendency rather than an absolute rule. In addition to differences in the speech signal, differences between the SR and the LR tone have also observed in EEG waveforms (Chen et al., 2022). The authors argue that different stages of encoding are involved in the production of Tone 3 sandhi.
@@
For a variety of reasons, the way that tones are realized in connected speech is different from the canonical form. Previous research has shown that tonal context is a crucial factor in tone (Xu, 1997). When placed in context, the realization of a given tone is greatly influenced by its preceding tone and following tone, resulting in tonal co-articulation. In addition to the influence of tonal context, speaking rate and the position of a word in the utterance also play a role (Wu et al., 2021; Yuan and Church, 2021). The faster a speaker talks, the less time is available for implementing changes in the muscles governing the vibration of the vocal cords, resulting in reduced pitch excursions (Cheng et al., 2013; Xu and Sun, 2002). Furthermore, a speaker’s speaking style has been found to be another factors predicting tonal variation, given that speakers has their own speaking (Stanford, 2016).
@@
The realization of tone is often also shaped by paralinguistic and sociolinguistic factors, including dialect, gender and a speaker’s emotional state. An analysis of the realization of tones in several automatic speech recognition corpora, registering the speech of some 2300 speakers from seven dialectal backgrounds (Tian et al., 2022), reported that tone sandhi was present not only for speakers of Beijing Mandarin, but also for speakers of other dialects of Mandarin, but with regional variation in the degree of assimilation.
@@
In addition, it has been demonstrated that gender plays a role in tonal realizations (Liang and Meng, 2011). Specifically, Tian et al. (2022) found that male speakers tend to produce a greater f0 rise than female speakers in producing both SR tone and LR tone. A speaker’s emotional state is also known to co-determine tonal realization (Chang et al., 2023; Zhang et al., 2006).
@@
Furthermore, T3-T3 tone sandhi has been argued to vary also with lexical frequency (Yuan and Chen, 2014; Tian et al., 2022). The study by Yuan and Chen (2014) found that the SR tone exhibited more f0 difference from the LR tone in more frequent words than in less frequent words. Replicating and extending this study, Tian et al. (2022) also reported a similar effect of word frequency across a range of dialects. One possible source for these frequency effects is that more frequent words tend to be realized with shorter spoken word duration, which in turn predicts reduced pitch excursion. A random forest analysis by Wu et al. (2023) shows that, among variation factors including prosodic position, word frequency, tonal contexts, and part of speech, lexical frequency has the most contribution to the SR tone’s pitch realization in naturally occurring journalist speech.
@@
Last but not least, some research has shown that fine phonetic detail may reflect subtle differences in meaning (Pierrehumbert et al., 2002). More recent studies supporting this possibility have investigated the spoken word duration of English heterographic homophones (Gahl and Baayen, 2022), the duration of English syllable-final /s/ (Tomaschek et al., 2021) and the duration of noun-verb conversion pairs (Lohmann, 2018). Chuang et al. (2024) report for conversational Taiwan Mandarin that the precise shape of the T2-T4 tone pattern is in part predictable from their meanings, while controlling for segment-related, speaker-related and contextual factors.
@@
In summary, incomplete neutralization, while commonly found for Beijing Mandarin, is however visible but not significant in Taiwan Mandarin. Tonal variation is co-determined by phonetic factors such as tonal context, speaking rate and speaker variability, by sociophonetic factors such as a speaker’s sociolect and gender, by the speaker’s emotional state, by frequency of use, and by the details of words’ meanings.
The present study is a follow-up of Yuan and Chen (2014), a study that addressed the acoustic characteristics of Tone 3 sandhi in spontaneous spoken Mandarin in large corpora of broadcast news and telephone conversations. The present study, however, investigates the tonal realization of disyllabic words with T2-T3 and T3-T3 in a corpus of face-to-face conversations in Taiwan Mandarin. Figure 1 provides examples of the tone contours of two words, 了解 (liao3jie3, to know) (upper panels) and 媒體 (mei2ti3, media) (lower panels) that illustrate the immense variation that characterizes pitch contours in this corpus. In our study, we leverage the power of the generalized additive model (GAM, Wood, 2017) to come grips with this variability and to come to a better understanding of four questions.
Figure 1: Example pitch contours for selected tokens of 了解 (liao3jie3, to know) (upper panels) and 媒體 (mei2ti3, media) (lower panels). Xu (1997) observed for laboratory speech that the f0 contours of T2-T3 and T3-T3 words consist of a slight fall, followed by a rise, and then a fall. A similar pattern is visible for the tokens at the left hand side, but very different realizations are found in the remaining panels.
@@
Third, does a word’s frequency of occurrence co-determine the shape of its pitch contour? Our hypothesis with respect to this question is that once word duration and speaking rate are taken into account, frequency does not further modulate pitch. This hypothesis is motivated by two considerations. First, frequency of use is a lexical property that is time-invariant, and it is unclear to us why it would modulate pitch over time. Second, if frequency of use would have a systematic effect on the realization of pitch, one would expect this modulation to be visible across all tonal patterns in the same way. No such effect has been reported in the existing literature.
Fourth, is the shape of a word’s pitch contour co-determined by its meaning? Based on the results obtained by Chuang et al. (2024) for the T2-T4 tone pattern in conversational Taiwan Mandarin, we expect that the fine details of how T2-T3 and T3-T3 tone contours are realized are word-specific and semantic in nature. If indeed words’ meanings co-determine their pitch contours, then this in turn raises the question of whether evidence for incomplete neutralization is robust when words and words’ senses are taken into account as control variables.

### v2

@@ ABSTRACT
In Standard Chinese, Tone 3 (the dipping tone) becomes Tone 2 (rising tone) when followed by another Tone 3. Previous studies have noted that this sandhi process may be incomplete, in the sense that the assimilated Tone 3 is still distinct from a true Tone 2. While Mandarin Tone 3 sandhi is widely studied using carefully controlled laboratory speech [57] and more formal registers of Beijing Mandarin [58], less is known about its realization in spontaneous speech, and about the effect of contextual factors on tonal realization. The present study investigates the pitch contours of two-character words with T2-T3 and T3-T3 tone patterns in spontaneous Taiwan Mandarin conversations. Our analysis makes use of the Generative Additive Mixed Model [[, GAMM,]]wood_generalized_2017 to examine fundamental frequency (F0) contours as a function of normalized time. We consider various factors known to influence pitch contours, including gender, duration, word position, bigram probability, neighboring tones, speaker, and also novel predictors, word and word sense [9]. Our analyses revealed that in spontaneous Taiwan Mandarin, T3-T3 words become indistinguishable from T2-T3 words, indicating complete sandhi, once the strong effect of word (or word sense) is taken into account.
KEYWORDS
@@
Tone 3 sandhi Tonal assimilation GAMM Word-specific tonal realization
@@
There is an increasing interest in spontaneous speech in phonetics, psycholinguistics, and related fields. While carefully controlled speech from laboratory experiments can provide valuable insights into specific aspects of speech production, spontaneous speech allows researchers to study authentic language use in everyday conversations. Spontaneous speech differs from laboratory speech in various aspects, such as the choice of words, syntactic patterns, tones and intonation, and phonological assimilation [47]. For instance, the pronunciation of Mandarin tones in spontaneous speech in everyday use differs from how it is described in textbooks and second-language learning classrooms, posing a huge challenge for second language learners of Mandarin.
@@
Mandarin Chinese is a tonal language, with each syllable being distinguished by a specific lexical tone in addition to vowels and consonants. This tonal differentiation contributes to distinguishing between words' meanings. The main acoustic correlate of Mandarin tones is the shape of words' fundamental frequency (F0) contours. Mandarin Chinese has four lexical tones: a high-level tone (T1), a rising tone (T2), a dipping tone (T3), and a falling tone (T4), along with a neutral or floating tone [6]. [6] describes the neutral tone as being unstressed, noticeably weaker in intensity, and shorter in duration. When a syllable is placed in a disyllabic word or connected speech, tone sandhi takes place. Tone sandhi refers to the modulation of a given syllable's lexical tones by the tones of other syllables in its context. Tone 3 sandhi is the most widely studied tone sandhi phenomenon in Mandarin Chinese. In a disyllabic word with the tone pattern T3-T3, the first T3 is typically pronounced as T2, leading to the perception that the T3-T3 tone pattern sounds similar to the T2-T3 tone pattern. There has been a long-standing debate about whether this tone sandhi results in incomplete or complete assimilation of initial T3 to an initial T2 [58].
@@
Although Tone 3 sandhi has long been studied for laboratory speech [42, 57], its realization in spontaneous speech has been less extensively investigated. It has been observed that in general, several standard acoustic cues for individual tones, such as F0, are greatly influenced by contextual factors or even reduced to a large degree in spontaneous speech, rendering them unreliable [3]. The majority of research in Mandarin Tone 3 sandhi is based on words that are articulated in isolation or words embedded in simple sentences or careful speech elicited in the lab using experimental tasks [49, 48]. In order to gain a more comprehensive understanding of the production of Tone 3 sandhi, it is essential to examine it in the speech style that speakers and listeners most frequently and commonly use in everyday settings. The study of [58], addresses the realization of Tone 3 sandhi as it is realized in spontaneous spoken Mandarin, using corpora of standard Mandarin telephone speech and Mandarin broadcast news speech.
The present study goes a step further, and investigates the realization of disyllabic words with T2-T3 and T3-T3 tone patterns in spontaneous spoken Taiwan Mandarin in face-to-face conversations. In what follows, section 1 provides further details on what is currently known about Tone 3 sandhi. Section 2 introduces our materials and the methods we used for data analysis. Results are presented in section 3. Our study concludes with a summary and discussion.
@@ 1 INCOMPLETE NEUTRALIZATION OF TONE 3 SANDHI
Most studies have reported incomplete neutralization for Tone 3 sandhi in Mandarin Chinese. The Sandhi Rising tone (SR tone, the first T3 in T3-T3) is found to exhibit lower pitch compared to the Lexical Rising tone (LR tone, T2 in T2-T3), with mean differences ranging from 3.2 Hz to 20 Hz [12, 26, 35, 40, 56, 57, 60]. The SR tone has also been described as tending to have less F0 excursion in the pitch contour across various varieties of Mandarin. For instance, [57] examined disyllabic non-word /ma-ma/ sequences with 16 possible Mandarin bi-tonal combinations embedded in carrier sentences produced by Beijing Mandarin native speakers. This study revealed that although both exhibited similar fall-rise-fall F0 contours, T3-T3 has a slightly lower pitch than T2-T3 throughout both syllables. This observation suggests for laboratory speech that the SR tone does not fully assimilate with the LR tone. However, most studies addressing Tone 3 sandhi in Taiwan Mandarin have reported that the SR tone and the LR tones are more similar to each other than in other varieties such as Beijing Mandarin. These differences in realization may remain visible but are not well supported by statistical analysis [8, 15, 35, 36].
@@
Only a few studies have examined Tone 3 sandhi in connected or spontaneous speech. An early study by [27] analyzed the speech of a speaker of Beijing Mandarin, who was asked to generate sentences using words presented in list format. Using discriminant analysis as classifier, he found that most of the tokens with the sandhi T3 were grouped with the lexical T3 rather than the lexical T2. This study concluded that speakers of Mandarin are adjusting the underlying T3, but the resulting SR tone is yet not equivalent to a T2.
@@
More recently, [58] analyzed the realization of Tone 3 sandhi in Mandarin bi-character words in connected speech from a large corpus of telephone conversations and formal news broadcasts. They found that, despite the remarkable similarity, the LR tone displays higher and longer F0 rising than the SR tone. This low-level acoustic difference in the magnitude and the time period of F0 rise is consistent with previous results obtained using carefully controlled speech. Following up on the study by [57], [54] analyzed Mandarin disyllabic words with 16 tonal combinations, but used large corpora of journalistic speech. They found that the SR tone in the T3-T3 sequence is often realized as other tones, suggesting that the tone sandhi rule in connected speech is more likely to be a tendency rather than an absolute rule. In addition to differences in the speech signal, differences between the SR and the LR tone have also observed in EEG waveforms [7]. The authors argue that different stages of encoding are involved in the production of Tone 3 sandhi.
@@
For a variety of reasons, the way that tones are realized in connected speech is different from the canonical form. Previous research has shown that tonal context is a crucial factor in shaping tone contours [57]. When placed in context, the realization of a given tone is greatly influenced by its preceding tone and following tone, resulting in tonal co-articulation. In addition to the influence of tonal context, speaking rate and the position of a word in the utterance also play a role [54, 59]. The faster a speaker talks, the less time is available for implementing changes in the muscles governing the vibration of the vocal cords, resulting in reduced pitch excursions [8, 55]. Furthermore, a speaker's speaking style has been found to be another factor predicting tonal variation, given that speakers has their own speaking [43].
@@
The realization of tone is often also shaped by paralinguistic and sociolinguistic factors, including dialect, gender and a speaker's emotional state. An analysis of the realization of tones in several automatic speech recognition corpora, registering the speech of some 2300 speakers from seven dialectal backgrounds [45], reported that tone sandhi was present not only for speakers of Beijing Mandarin, but also for speakers of other dialects of Mandarin, but with regional variation in the degree of assimilation.
@@
In addition, it has been demonstrated that gender plays a role in tonal realizations [29]. Specifically, [45] found that male speakers tend to produce a greater F0 rise than female speakers in producing both SR tone and LR tone. A speaker's emotional state is also known to co-determine tonal realization [5, 61].
@@
Furthermore, T3-T3 tone sandhi has been argued to vary also with lexical frequency [58, 45]. The study by [58] found that the SR tone exhibited more F0 difference from the LR tone in more frequent words than in less frequent words. Replicating and extending this study, [45] also reported a similar effect of word frequency across a range of dialects. One possible source for these frequency effects is that more frequent words tend to be realized with shorter spoken word duration, which in turn predicts reduced pitch excursion. A random forest analysis by [53] shows that, among varying factors including prosodic position, word frequency, tonal contexts, and part of speech, lexical frequency has the most contribution to the SR tone's pitch realization in naturally occurring journalist speech.
@@
Last but not least, some research has shown that fine phonetic detail may reflect subtle differences in meaning [37]. More recent studies supporting this possibility have investigated the spoken word duration of English heterographic homophones [17], the duration of English syllable-final /s/ [46] and the duration of noun-verb conversion pairs [30]. [9] report for conversational Taiwan Mandarin that the precise shape of the T2-T4 tone pattern is in part predictable from their meanings, while controlling for segment-related, speaker-related and contextual factors.
@@
In summary, incomplete neutralization, while commonly found for Beijing Mandarin, is however visible but not significant in Taiwan Mandarin. Tonal variation in general is co-determined by phonetic factors such as tonal context, speaking rate and speaker variability, by sociophonetic factors such as a speaker's sociolect and gender, by the speaker's emotional state, by frequency of use, and by the details of words' meanings.
The present study is a follow-up of [58], a study that addressed the acoustic characteristics of Tone 3 sandhi in spontaneous spoken Mandarin in large corpora of broadcast news and telephone conversations. The present study, however, investigates the tonal realization of disyllabic words with T2-T3 and T3-T3 in a corpus of face-to-face conversations in Taiwan Mandarin. Figure 1 provides examples of the tone contours of two words, 了解 (liao3jie3, to know) (upper panels) and 媒体 (mei2ti3, media) (lower panels) that illustrate the immense variation that characterizes pitch contours in this corpus. [57] observed for laboratory speech that the F0 contours of T2-T3 and T3-T3 words consist of a slight fall, followed by a rise, and then a fall. A similar pattern is visible for the tokens at the left hand side (ZWH_GY_1763_了解 and LJS_GY_8814_媒体), but very different realizations are found in the remaining panels. The pitch contour of ZWH_GY_1763_了解, which has a relatively long duration (0.40 s), generally follows the pattern described in [57], except that the F0 of the initial consonant [tC] of 解 is not present. In contrast, LJS_GY_6870_媒体 with a short duration (0.26 s), exhibits a reduced pitch contour that appears as a falling pattern. Besides, tonal co-articulation is another important factor of tonal variations. For example, the F0 contour of the token GYX_GY_4564_了解 is high and rising at the end, rather than expected low and falling. This is because it occurs in an utterance with exclamatory intonation 太了解啦 (tai4liao3jie3la0, knows very well!) and is followed by 啦 (la0), which in this utterance is realized with relatively high pitch. The token KCZX_GY_1537_了解 exhibits an initially level pitch contour because it is preceded by 我 (wo3, I), whose Tone 3 is often realized as a low tone in Taiwan Mandarin [15, 14].
Figure 1: Example pitch contours for selected tokens of 了解 (liao3jie3, `to know') (upper panels) and 媒体 (mei2ti3, `media') (lower panels). The pitch contours are color-coded by gender, with orange indicating female speakers and blue indicating male speakers. The vertical dashed line represents token-specific syllable boundary on the normalized time scale. The duration of each token (in seconds) is noted in the bottom of each panel.
In our study, we leverage the power of the generalized additive model [[, GAM,]]wood_generalized_2017 to come to grips with this variability and to come to a better understanding of four questions.
@@
Third, does a word's frequency of occurrence co-determine the shape of its pitch contour? Our hypothesis with respect to this question is that once word duration and speaking rate are taken into account, frequency does not further modulate pitch. This hypothesis is motivated by two considerations. First, frequency of use is a lexical property that is time-invariant, and it is unclear to us why it would modulate pitch over time. Second, if frequency of use would have a systematic effect on the realization of pitch, one would expect this modulation to be visible across all tonal patterns in the same way. No such effect has been reported in the existing literature.
Fourth, is the shape of a word's pitch contour co-determined by its meaning? Based on the results obtained by [9] for the T2-T4 tone pattern in conversational Taiwan Mandarin, we expect that the fine details of how T2-T3 and T3-T3 tone contours are realized are word-specific and semantic in nature. If indeed words' meanings co-determine their pitch contours, then this in turn raises the question of whether evidence for incomplete neutralization is robust when words and words' senses are taken into account as control variables.
In the current study, the statistical analysis will be conducted on entire words, including consonant onsets and vowels. For example, for 媒体 (mei2ti3), the analysis is performed on [meIthi]. This approach aligns with the general framework of the Discriminative Lexicon Model [9, 21], which focuses on the relationship between semantics and phonetic realization. Previous work has shown that semantics, operationalized using contextualized embeddings, predicts the pitch contours of word tokens [9, 31]. Therefore, our focus is on how pitch is realized on whole onomasiological words rather than on their constituent morphemes or syllables.